Tom Smith purchases 100 shares of DOUBLE Systems stock for $63 per share and wishes to...

  1. Tom Smith purchases 100 shares of DOUBLE Systems stock for $63 per share and wishes to hedge his position by writing a 100 share call option on his holdings. The option per share has a $65 strike price and a premium of $8.75. If the stock is selling at $64 at the time of expiration, what will be the overall dollar gain or loss on this covered option play? (Consider the change in stock value as well as the gain or loss on the option.)

    1. $975.00

    2. $875.00

    3. $775.00

    4. $100.00

    5. $87.50
Answer #1

The strategy is called covered call. It consists of a long stock and a short call.

Stock price at expiration = 64

Number of stock = 100

Buying price of stock = 63

profit from stock at expiration = (64 - 63) × 100 = $100

Premium from options = 8.75 × 100 = $875

The option is not exercised as the strike price ($65) is more than the stock price.

Therefore total profit = profit from stock + option premium

= $100 + $875

= $975